FriskIsNice:

Is the square root of 66 a rational or irrational number?

Extrinix:

\(\sqrt{66}\) Rational: is a fraction Irrational: is not a fraction So from this, what do you think the answer is?

FriskIsNice:

@extrinix wrote:
\(\sqrt{66}\) Rational: is a fraction Irrational: is not a fraction So from this, what do you think the answer is?
It’s irrational right? Cause it’s not a fraction, and it dose t repeat.

Extrinix:

It's looking for an answer of \(\sqrt{66}\), so no.